# Democratizing AI scientists using ToolUniverse

### v1

@@ 1 Overview of ToolUniverse
Core Components. ToolUniverse include Tool Discover and Tool Manager for discovering and integrating tools, a Tool Finder to search for options from over 600 candidates based on user requirements, and a Tool Caller for execution. For complex tasks, the Tool Composer assembles multiple tools into a composite workflow. The Tool Optimizer utilizes a built-in multi-agent system to refine tool specifications, ensuring they better align with the tool’s actual behavior. Leveraging these components and the unified AI-tool interaction protocol, ToolUniverse empowers creating AI scientists: specialized AI models that combine reasoning with access to curated tools to perform complex research tasks.
@@ 3 Core Components of ToolUniverse
ToolUniverse operates through a set of core components designed for comprehensive tool management. Its capabilities include Tool Discover to discover new tools and Tool Manager to integrate tools into ToolUniverse, a Tool Finder to search for suitable options from over 600 candidates based on user requirements, and a Tool Caller for execution. For complex tasks, the Tool Composer assembles multiple tools into a new, composite workflow. The Tool Optimizer utilizes a built-in multi-agent system to refine tool specifications, ensuring they better align with the tool’s actual behavior.
@@ 3.5 Tool Optimizer
3.6 Tool Discover
Tool Discover automatically generates new tools, including both specifications and executable code, from high-level natural language descriptions. Leveraging the workflow composition capabilities of ToolUniverse via Tool Composer, we construct an agentic multi-stage pipeline that transforms a plain-text functional request into a production-ready tool with minimal human intervention. The process integrates tool discovery, structured specification generation, code implementation, and iterative quality refinement, ensuring that the final tool adheres to functional requirements and ecosystem conventions.
@@ 3.6 Tool Discover
Core principles and architecture. The Tool Discover system is built around four core principles: pattern-guided generation that reuses functional patterns and conventions from existing tools to ensure ecosystem consistency; structured specification synthesis that transforms unstructured requests into tool specifications; automated code generation and validation that produces executable implementations with integrated testing; and iterative refinement that uses feedback from analysis and testing to drive targeted improvements.
@@ 4.1 Building an AI Scientist from an LLM or an LRM
Figure 2(b) illustrates an example of building an AI scientist using an LLM, such as Claude, together with ToolUniverse. The process involves only three steps.
@@ 4.3 Building an AI Scientist from a Specialized AI Agent
ML models Tools that apply machine learning algorithms to tasks like prediction, classification, or generation.
AI agents Tools that operate autonomously to perceive environments, make decisions, and take actions toward goals.
Scientific software packages Biomedical software packages engineered to facilitate diverse scientific tasks, experiments, and data analysis workflows.
Human expert feedback Tools that incorporate evaluations or input from human experts into automated or semi-automated processes.
Robotics Tools involving physical or simulated machines capable of sensing, reasoning, and acting in the world.
Databases Tools that store, manage, and query structured or semi-structured data efficiently, including relational, tabular, and hierarchical databases.
Embedding stores Tools that store and retrieve vectorized representations of data for use in machine learning tasks.
APIs Interface tools that allow different software systems or tools to communicate and exchange data.
Supplementary Table 2: Number of tools and example tools for each category in ToolUniverse.
Tool category Number Example of tools in ToolUniverse
@@
AI agents 38 HypothesisGenerator, ExperimentalDesignScorer, MedicalLiteratureReviewer
Scientific software packages 164 get_biopython_info, get_pyscreener_info, get_pykalman_info
Human expert feedback 6 consult_human_expert, get_expert_response, get_expert_status
Robotics 1 communicate_with_ros_robot
Databases 84 drugbank_get_drug_pathways_and_reactions_by_name, HPA_get_comprehensive_gene_details_by_ensembl_id, drugbank_full_search
Embedding stores 4 embedding_tool_finder, embedding_database_search, embedding_database_add
APIs 281 FDA_get_active_ingredient_info_by_drug_name, PubChem_search_compounds_by_substructure, OpenTargets_get_associated_targets_by_disease_efoId
@@ 6 Evaluation of Tools
  • Input–Output Sampling. For each tool, we generated diverse sample inputs and recorded corresponding outputs. These samples were constructed to cover typical use cases as well as edge conditions, ensuring broad coverage of tool functionality.
@@
  • Regular Maintenance and Bug Reporting. To ensure continued reliability of tools within ToolUniverse, we perform regular maintenance and monitoring. A structured bug reporting system allows issues to be identified by users and promptly addressed by the ToolUniverse team. These practices provide ongoing quality assurance and safeguard the long-term usability of the ecosystem.
Beyond these internal steps, ToolUniverse prioritizes tools from trusted sources. Many included tools have been published, peer-reviewed, and verified through prior use by established scientific communities, such as the NIH, FDA, and other regulatory or research agencies. By relying on previously validated resources, we reduce the risk of introducing spurious or unverified functionality into the ecosystem. These evaluation measures ensure that tools incorporated into ToolUniverse meet standards of correctness, reproducibility, and scientific reliability, providing a foundation for dependable AI-assisted discovery.
@@ 7 Further Details on the Case Study of ToolUniverse
Finally, we use the AI scientist’s outputs to select two structural analogs for further evaluation: CHEM-BL2347006 and CHEMBL3970138. These analogs are in fact the same molecule despite slight differences in their SMILE representation. This molecule places within the top five candidates for highest probability of binding the target (shown in Supplementary Table LABEL:table:chembl_values), has the strongest predicted binding affinity, and places within the top five for lowest predicted BBB penetrance.
@@
This case study demonstrates how ToolUniverse can be used to conduct flexible, multi-domain scientific research. It breaks out of the paradigm of a narrowly-focused bioinformatics pipeline; instead, the AI scientist leverages tools and data modalities across chemistry, human biology, and patent literature to retrieve multiple lines of evidence in supporting its final drug candidates. Moreover, by recovering pravastatin, a statin with better off-target binding properties compared to the starting compound of lovastatin, the case study demonstrates the ability of ToolUniverse-mediated research to recover—and therefore advance—current science.

### v2

@@ 1 Overview of ToolUniverse
Core Components. ToolUniverse include Tool Discoverer and Tool Manager for discovering and integrating tools, a Tool Finder to search for options from over 600 candidates based on user requirements, and a Tool Caller for execution. For complex tasks, the Tool Composer assembles multiple tools into a composite workflow. The Tool Optimizer utilizes a built-in multi-agent system to refine tool specifications, ensuring they better align with the tool’s actual behavior. Leveraging these components and the unified AI-tool interaction protocol, ToolUniverse empowers creating AI scientists: specialized AI models that combine reasoning with access to curated tools to perform complex research tasks.
@@ 3 Core Components of ToolUniverse
ToolUniverse operates through a set of core components designed for comprehensive tool management. Its capabilities include Tool Discoverer to discover new tools and Tool Manager to integrate tools into ToolUniverse, a Tool Finder to search for suitable options from over 600 candidates based on user requirements, and a Tool Caller for execution. For complex tasks, the Tool Composer assembles multiple tools into a new, composite workflow. The Tool Optimizer utilizes a built-in multi-agent system to refine tool specifications, ensuring they better align with the tool’s actual behavior.
@@ 3.5 Tool Optimizer
3.6 Tool Discoverer
Tool Discoverer automatically generates new tools, including both specifications and executable code, from high-level natural language descriptions. Leveraging the workflow composition capabilities of ToolUniverse via Tool Composer, we construct an agentic multi-stage pipeline that transforms a plain-text functional request into a production-ready tool with minimal human intervention. The process integrates tool discovery, structured specification generation, code implementation, and iterative quality refinement, ensuring that the final tool adheres to functional requirements and ecosystem conventions.
@@ 3.6 Tool Discoverer
Core principles and architecture. The Tool Discoverer system is built around four core principles: pattern-guided generation that reuses functional patterns and conventions from existing tools to ensure ecosystem consistency; structured specification synthesis that transforms unstructured requests into tool specifications; automated code generation and validation that produces executable implementations with integrated testing; and iterative refinement that uses feedback from analysis and testing to drive targeted improvements.
@@ 4.1 Building an AI Scientist from an LLM or an LRM
Figure 2b illustrates an example of building an AI scientist using an LLM, such as Claude, together with ToolUniverse. The process involves only three steps.
@@ 4.3 Building an AI Scientist from a Specialized AI Agent
ML models Tools that apply machine learning algorithms to tasks like prediction, classification, or generation. Examples include foundation models, language models, and reasoning models.
Agents Tools that operate autonomously to perceive environments, make decisions, and take actions toward goals. Examples include research planning agents, hypothesis generation agents, and data analysis agents.
Scientific packages Software packages engineered to facilitate diverse scientific tasks, experiments, and data analysis workflows. Examples include computation packages, biomedical libraries, and scientific simulators.
Automation Tools involving physical or simulated machines capable of sensing, reasoning, and acting in the world. Examples include lab automation tools and instrument control policies.
Workflows Tools that enable complex, multi-step scientific workflows. Examples include orchestration schedulers and RAG pipelines.
Visualization Tools that facilitate the display and communication of scientific data and results. Examples include dashboards and plot and chart generators.
Feedback and safety Tools that incorporate evaluations or input from human experts to ensure safety and goal alignment. Examples include privacy guardrails, security checklists, and human-in-the-loop tools.
Local and remote datasets Tools that store, manage, and query structured or semi-structured data efficiently, including relational, tabular, and hierarchical datasets. Examples include knowledge bases and typed API clients.
Experimentation Tools that support the design and management of experiments. Examples include protocol generators and LIMS integration tools.
Search and embeddings Tools that store and retrieve vectorized representations of data for use in machine learning tasks. Examples include vector search tools and embedding generators.
Supplementary Table 2: Number of tools and example tools for each category in ToolUniverse. The table shows the number of tools as of October 11, 2025. Note that ToolUniverse is continually expanded with new tools.
Tool category Number Examples of tools in ToolUniverse
@@
Agents 56 HypothesisGenerator, CodeQualityAnalyzer, MedicalLiteratureReviewer
Scientific packages 164 get_biopython_info, get_pyscreener_info, get_pykalman_info
Automation 8 communicate_with_ros_robot, mcp_auto_loader, mcp_client
Workflows 11 biomarker_discovery_workflow tool_optimizer tool_discoverer
Visualization 3 visualize_protein_structure_3d visualize_molecule_2d visualize_molecule_3d
Feedback and safety 6 consult_human_expert, get_expert_response, get_expert_status
Local and remote datasets 391 drugbank_get_drug_pathways_and_reactions_by_name, HPA_get_comprehensive_gene_details_by_ensembl_id, FDA_get_active_ingredient_info_by_drug_name, OpenTargets_get_associated_targets_by_disease_efoId
Search and embeddings 4 embedding_tool_finder, embedding_database_search, embedding_database_add
Experimentation 3 experimental__design_scorer, protocol_optimizer, extract_clinical_trial_outcomes
@@ 6 Evaluation of Tools
  • Input-output sampling. For each tool, we generated diverse sample inputs and recorded corresponding outputs. These samples were constructed to cover typical use cases as well as edge conditions, ensuring broad coverage of tool functionality.
@@
  • Regular maintenance and bug reporting. For continued reliability of tools within ToolUniverse, we perform regular maintenance and monitoring. A structured bug reporting system allows issues to be identified by users and promptly addressed by the ToolUniverse team. These practices provide ongoing quality assurance and safeguard the long-term usability of the ecosystem.
Beyond these processes, ToolUniverse prioritizes tools from trusted sources. Many included tools have been published, peer-reviewed, and verified through prior use by established scientific communities, such as the NIH, FDA, and other regulatory or research agencies. By relying on previously validated resources, we reduce the risk of introducing spurious or unverified functionality into the ecosystem. These evaluation measures ensure that tools incorporated into ToolUniverse meet standards of correctness, reproducibility, and scientific reliability, providing a foundation for dependable AI-assisted discovery.
@@ 7 Further Details on the Case Study of ToolUniverse
Finally, we use the AI scientist’s outputs to select two structural analogs for further evaluation: CHEMBL2347006 and CHEMBL3970138. These analogs are in fact the same molecule despite slight differences in their SMILES representation. This molecule places within the top five candidates for highest probability of binding the target (shown in Supplementary Table LABEL:table:chembl_values), has the strongest predicted binding affinity, and places within the top five for lowest predicted BBB penetrance.
@@
This case study demonstrates how ToolUniverse can be used to conduct flexible, multi-domain scientific research. It breaks out of the paradigm of a narrowly-focused bioinformatics pipeline; instead, the AI scientist leverages tools and data modalities across chemistry, human biology, and patent literature to retrieve multiple lines of evidence in supporting its final drug candidates.